OpenStudy (anonymous):

angela ate at the same restaurant four times. each time she ordered a salad and left a $5 tip. she spent a total of $54. write and solve an equation to find the cost of each salad.

OpenStudy (anonymous):

x= price of salad Each visit is (x+5). She visited 4 times, so... 4(x+5)=54 4x+20=54 4x=34 x=$8.50

OpenStudy (anonymous):

$7 per salad

OpenStudy (anonymous):

54-(4 x 4.5)÷4 54- 26= 28 28÷4=7 $7 per salad

OpenStudy (anonymous):

Where did you get the 4.5? I'm not seeing the logic in your math. Where are the variables?

OpenStudy (anonymous):

where did u get the five

OpenStudy (anonymous):

The 5 came from the $5 tip left each time.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

You're welcome!